40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4367     Entry time: Wed Mar 2 16:51:53 2011
Author: steve 
Type: Configuration 
Category: Green Locking 
Subject: mech shutter in place at the south end 

I moved old POX shutter from ITMY optical table to the south end. MEDM POX mechanical shutter screen is now closing the green beam  injection into the Y arm.

I kluged in a 40m long bnc cable that Alberto left on the floor for control. It is labelled POX-sht  This is a temporary set up.

ELOG V3.1.3-